OpenStudy (anonymous):

Need help with system: 5x+4y=-30 3x-9y=-18

OpenStudy (anonymous):

In the second equation you should first divide each term by three to obtain x-3y=-6 and then solve for x and get x= 3y-6. Then using substitution you'll get 5(3y-6)+4y=-30. Then distribute the 5 to get, 15y-30+4y=-30. By combining like terms, you will get 19y=0 so y=0. You already solved for x in terms of y, so since x=3y-6 and y=0 so x=-6.